A firm manufactures snow blowers. The demand for snow blowers follows a seasonal pattern with highs during the winter months and lows during the summer months. The firm’s cost data and sales forecasts are as given below.

Quarter

Sales Forecast (units)

Spring

80,000

Summer

50,000

Fall

120,000

Winter

150,000

Hiring cost is $100 per worker

Firing cost is $500 per worker

Inventory carrying cost is $0.50 per unit per quarter

Production per employee is 1,000 units per quarter

Beginning work force is 100 workers.  

a) Calculate a level plan

b) Calculate a chase plan

c) Which plan (level or chase) is economical ?
Answer #1
Quarter Sales Forecast (units) Production Workers needed Hiring firing beginning inventory Ending inventory
Spring 80,000 100000 100 0 0 0 20000
Summer 50,000 100000 100 0 0 20000 70000
Fall 120,000 100000 100 0 0 70000 50000
Winter 150,000 100000 100 0 0 50000 0

cost of level strategy

= production cost + inventory cost = PC+ avg inventory x0.5 =PC+140000x0.5

= PC+70000

2. Chase strategy

Quarter Sales Forecast (units) Production Workers needed Hiring firing
Spring 80,000 80000 80 0 20
Summer 50,000 50000 50 0 30
Fall 120,000 120000 120 70 0
Winter 150,000 150000 150 30 0

cost of this strategy = Production cost + hiring cost + firing cost

= PC + 100x100 +50x500

= PC + 35000

Chase plan is more economical by 35000.